# SCSA: Exploring the Synergistic Effects Between Spatial and Channel Attention

### v1

@@ E-A Visualization of ERFs
E-B Computational Complexity
Given an input ${X\in\mathbb{R}^{C\times H\times W}}$, a pooling size of $P\times P$, and a depth-wise convolutional kernel size of $K\times K$, we sequentially consider the impact of dimension decoupling, depth-shared 1D convolutions, normalization, progressive compression, and channel-wise self-attention, which collectively constitute the SCSA module. For simplicity of observation, we ignore the coefficients. The computational complexities of SCSA are:
@@ E-B Computational Complexity
E-C Inference Throughput Evaluation
@@
As demonstrated in Tables III and IV, we evaluate the throughput of SCSA’s individual components in ablation experiments and compare the throughput across various benchmark models using different attention mechanisms. We obtain results using an GeForce RTX 4090 GPU at 224x224 resolution. Specifically, As illustrated in Table III, although SCSA is slightly slower than pure channel attention, it outperforms most hybrid attention mechanisms, including CBAM, SANet, EMA, and CPCA, and achieves the highest accuracy. Table III indicates our design optimizes balance in model complexity, inference speed, and accuracy.
Appendix F Limitations
@@
We demonstrated that our SCSA, a plug-and-play synergistic attention method, excels in image classification, object detection, and instance and semantic segmentation tasks. Although we are committed to exploring the synergistic effects across various dimensions and have empirically validated the effectiveness of leveraging multi-semantic spatial information to guide channel recalibration and enhance feature interactions for mitigating semantic differences, inference latency remains a significant challenge in real-world deployment. Our approach achieves an optimal balance of model parameters, accuracy, and inference speed at an appropriate model width. However, at larger widths, the primary bottleneck in inference speed is the use of depth-wise convolutions within the construction of a mutli-semantic spatial structure, which have low FLOPS, frequently access memory, and exhibit low computational density[26]. We believe that the positioning and quantity of attention modules should be optimized based on specific tasks and scenarios to ensure peak performance. In the future, we will explore attention synergy across multiple dimensions to ensure that attention mechanisms across different dimensions complement and enhance each other.

### v2

@@ 4.5 Ablation study
Branch.
The richness of semantic feature capture in SMSA is determined by the number of branches and the convolution kernel sizes used in each branch. Each branch is designed to learn a distinct sub-features. Reducing the number of branches weakens the module’s ability to extract inherent multi-semantic features. To assess the impact of capturing different semantic features on model performance, we conducted experiments with varying branch numbers and convolution kernel sizes. As shown in the "Branch" section of table 1, the accuracy of the dual-branch structure surpasses that of the single-branch, while the four-branch structure further outperforms the dual-branch. This results supports the effectiveness of our multi-branch, multi-scale structure in capturing diverse semantic patterns across sub-features, thereby enhancing the model’s representational capacity. As the number of branches increases, the model’s memory access overhead also rises, resulting in a decrease in inference speed.
5 Visualization and Analysis
5.1 Visualization of Attention
We evaluate the effectiveness of our method in mitigating semantic disparities and enhancing consistency by ensuring appropriate attention to key regions. As shown in fig. 4, compared to other state-of-the-art attention mechanisms, our SCSA distinctly focuses on multiple key regions, significantly reducing critical information loss while providing rich feature information. We also visualize the components of SCSA, including the SMSA and PCSA modules. In the absence of the PCSA module to address semantic disparities, the distribution of activation intensity remains inadequately balanced. Without the SMSA module for guiding multi-semantic spaces, the focus on important regions may be limited.
5.2 Visualization of Effective Receptive Field
As depicted in fig. 5, leveraging the spatial structure of multi-semantic modeling, our SCSA has achieved a broader perceptual area. A larger effective receptive field (ERF) is beneficial for the network to utilize rich contextual information for collective decision-making, which is one of the important factors for performance improvement. To verify that the performance of our method benefits from a larger ERF, we randomly sample 300 images of different categories from the ImageNet-1K validation set [44], measure the contribution of each pixel on the original image to the center point of the output feature maps of the third and fourth stages of the model, and quantify the range of the ERF with the gradient values weighted and normalized. The visualization results demonstrate that as the network layers deepen, the ERF of our SCSA becomes increasingly evident, confirming our hypothesis and the effectiveness of our method.
5.3 Computational Complexity
Given an input ${X\in\mathbb{R}^{B\times C\times H\times W}}$, a pooling size of $P\times P$, and a depth-wise convolutional kernel size of $K\times K$, we sequentially consider the impact of dimension decoupling, depth-shared 1D convolutions, normalization, progressive compression, and channel-wise self-attention, which collectively constitute the SCSA module. For simplicity of observation, we ignore the coefficients. The computational complexities of SCSA are:
@@ 5.3 Computational Complexity
Figure 6: Detection results are visualized on Faster R-CNN [43], Cascade R-CNN [3], and RetinaNet [29] by respectively selecting two random samples from MSCOCO validation set [30] and comparing our SCSA with the ResNet-50 [19] baseline to demonstrate the effectiveness of our method.
5.4 Inference Throughput Evaluation
@@
As demonstrated in table 1 and table 2, we evaluate the throughput of SCSA’s individual components in ablation experiments and compare the throughput across various benchmark models using different attention mechanisms. We conduct our experiments using a GeForce RTX 4090 GPU at a resolution of 224x224, with a batch size of 32 to simulate real-world applications and maximize GPU utilization. To minimize variability, we repeat 100 times for each attention mechanism and report the average inference time. Specifically, As illustrated in table 2, although SCSA is slightly slower than pure channel attention, it outperforms most hybrid attention mechanisms, including CBAM, SANet, EMA, and CPCA, and achieves the highest accuracy.
5.5 Qualitative Results of Object Detection
As shown in fig. 6, our method demonstrates superior performance in challenging scenarios, including obstruction, dense environments, clusters of small objects, and low-light conditions.
Figure 7: Visualization of instance segmentation results using the Mask R-CNN [18]. Each instance depicted in a distinct color.
5.6 Qualitative Results of Instance Segmentation
As shown in fig. 7, our method segments obscured and overlapping objects more comprehensively and accurately, achieving higher confidence scores. These results underscore the benefits of our method in leveraging multi-semantic information to better perceive the contextual space of relevant objects.
Figure 8: Visualization of semantic segmentation results using the UperNet [59] on ADE20K [67].
5.7 Qualitative Results of Semantic Segmentation
It can be observed from fig. 8 that our method significantly improves the segmentation of objects that overlap and are semantically adjacent, effectively distinguishing between scenarios such as spectators seated on chairs and toilets near bathtubs.
6 Limitations
@@
We demonstrated that our SCSA, a plug-and-play synergistic attention method, excels in image classification, object detection, and instance and semantic segmentation. Although we are committed to exploring the synergistic effects across various dimensions and have empirically validated the effectiveness of leveraging multi-semantic spatial information to guide channel recalibration and enhance feature interactions for mitigating semantic disparities, inference latency remains a significant challenge in real-world deployment. Our approach achieves an optimal balance of model parameters, accuracy, and inference speed at an appropriate model width. However, at larger widths, the primary bottleneck in inference speed is the use of depth-wise convolutions and branching within the construction of a mutli-semantic spatial structure, which have low FLOPS, frequently access memory, and exhibit low computational density [4, 51, 36]. We believe that the positioning and quantity of these plug-and-play attention modules should be optimized based on specific tasks and scenarios to ensure peak performance. In the future, we will investigate more lightweight and faster plug-and-play attention mechanisms, exploring the synergistic relationships across different dimensions.
7 Conclusion
In this study, we analyze the limitations of most plug-and-play attention methods in leveraging the inherent multi-semantic information of features across spatial and channel dimensions, as well as the challenges posed by semantic disparities. To address these issues, we propose a novel plug-and-play Spatial and Channel Synergistic Attention (SCSA) mechanism, which incorporates dimension decoupling, lightweight multi-semantic guidance, and semantic disparity mitigation. SCSA leverages multi-semantic spatial attention to guide the learning of diverse channel features, followed by single-head self-attention in the channel dimension to alleviate semantic disparities and promote semantic interaction. Extensive experiments demonstrate that SCSA consistently outperforms state-of-the-art attention mechanisms on widely used benchmarks, showing enhanced performance and robust generalization capabilities. We hope our work encourages further exploration of synergistic properties across multiple dimensions in various domains.